OpenStudy (anonymous):

A local hamburger shop sold a combined total of 595 hamburgers and cheeseburgers on Tuesday. There were 55 fewer cheeseburgers sold than hamburgers. How many hamburgers were sold on Tuesday?

OpenStudy (zpupster):

x+y = 595 cb and hb total sold then cb x+55=y cb+55 equals total hb sold substitute x+x+55=595 solve for x

OpenStudy (anonymous):

thank you zpupster